OpenStudy (anonymous):

What is the equation in point-slope form of the line passing through (0, 5) and (−2, 11)?

OpenStudy (anonymous):

Firstly find the slope : \[Slope = \frac{y_2 - y_1}{x_2 - x_1}\]

OpenStudy (anonymous):

\(x_1 = 0,\ y_1 = 5,\ x_2 = -2, \ y_2 = 11\)

OpenStudy (anonymous):

-3

OpenStudy (anonymous):

So, it looks good to me.. :)

OpenStudy (anonymous):

\[Slope = \frac{11-5}{-2} = -3\]

OpenStudy (anonymous):

Now, equation in point-slope form is given as : \(y = mx + b\), where \(m\) is the slope and \(b\) is y-intercept, Right?

OpenStudy (anonymous):

\(m\) we have found as \(-3\) just now.. So, let us put it in that equation.

OpenStudy (anonymous):

\(y = -3x + b\) Now, we need to find \(b\)..

OpenStudy (anonymous):

Just substitute any one point in that equation, because, if line passes through \((0,5)\), then this point must satisfy the equation : So put : \(x=0, \ y = 5\) in that equation and find \(b\)..

OpenStudy (anonymous):

When you put \(x=0\) and \(y=5\), in \(y = -3x + b\), then what the value of \(b\) you will get?

OpenStudy (anonymous):

8?